\begin{thebibliography}{9990}

\bibitem[AC]{AC}
Ablowitz M., Clarkson P., \textit{Solitons, Nonlinear Evolution Equations and Inverse Scattering}, Cambridge University Press, Cambridge, 1991

\bibitem[A1]{A1}
Andreev V.A., Supersymmetric generalized Toda lattice. In: M.~Markov,
V.~Man'ko (eds.) \textit{Proc. International Conf. Group-theoretical
Methods in Physics},  Vol. I (Yurmala, 1985), VNU Sci. Press, Utrecht, 1986, 315--321

\bibitem[A2]{A2}
Andreev V.A., Odd bases of Lie superalgebras and integrable equations. Theoret. and Math. Physics, 72:1 (1987) 758--764

%\bibitem[AB]{AB}
%Andreev V.A., Burova M.V., Lower Korteweg--de Vries equations and supersymmetric structure of the sine-Gordon and Liouville equations. Theoret. and Math. Phys., 85:3 (1990), 1275--1283

\bibitem[AN]{AN}
Arnold V. I., Novikov S. P. (eds.) Dynamical Systems VII.
Integrable Systems Nonholonomic Dynamical Systems. Springer-Verlag Berlin Heidelberg (1994) vii+344 pp.

\bibitem[BL]{BL}
Bernstein J.N., Leites D., Integral forms and Stokes's formula on
supermanifolds. Funct. Anal. Appl.,
\textbf{11} (1977), no.~1, 55--56

%\bibitem[BGL]{BGL}
%Bouarroudj S., Grozman P., Leites D., Classification of finite-dimensional modular Lie superalgebras with indecomposable Cartan
%matrix. Symmetry, Integrability and Geometry: Methods and
%Applications (SIGMA), v.~5 (2009), 060, 63 pages;
%\url{https://arxiv.org/pdf/math.RT/0710.5149}

\bibitem[BGLS]{BGLS}
Bouarroudj S., Grozman P., Leites D., Shchepochkina I., Minkowski superspaces and superstrings as almost
real-complex supermanifolds.  Theor. and Mathem. Physics, Vol. 173, no.~3,
(2012) 1687--1708; \url{https://arxiv.org/pdf/1010.4480}

\bibitem[BLS]{BLS}
Bouarroudj S., Leites D., Shang J., Computer-aided study of double extensions of restricted Lie superalgebras preserving the non-degenerate closed \protect{2-forms} in characteristic 2. Experimental Math. \url{https://doi.org/10.1080/10586458.2019.1683102};
\url{https://arxiv.org/pdf/1904.09579}


%\bibitem[BLLS]{BLLS}
%Bouarroudj S., Leites D., Lozhechnyk O., Shang J., The roots of exceptional modular Lie superalgebras with Cartan matrix. Arnold Math. J. (2020) \url{https://doi.org/10.1007/s40598-020-00135-x}; \url{https://arxiv.org/pdf/1904.09578}

%\bibitem[BCDS]{BCDS}
%Braden H. W., Corrigan E., Dorey P. E., Sasaki R., Affine Toda field theory and exact S-matrices. Nuclear Phys. B 338 (1990), no. 3, 689--746.



\bibitem[CCLL]{CCLL}
Chapovalov D., Chapovalov M., Lebedev A.,
Leites D., The classification of almost affine (hyperbolic) Lie
superalgebras. J. Nonlinear Math. Phys., v.~17 (2010), Special issue
1, 103--161; \url{https://arxiv.org/pdf/0906.1860}

%\bibitem[CTW]{CTW}
%Coldea R., Tennant D.A., Wheeler E.M., Wawrzynska E., Prabhakaran, Telling M., Habnicht D.K., Smeibidl P., and Kiefer K., Quantum criticality in an Ising chain: Experimental evidence for emergent $E_8$ symmetry, Science, 327 (2010), Issue 5962, 177--180; \url{https://arxiv.org/pdf/1103.3694} (cond-mat)

%\bibitem[dSKV]{dSKV}
%De Sole A., Kac V. G., Valeri D.,  On Lax operators. Japan. J. Math. 17 (2022) 63--116; \url{https://arxiv.org/pdf/2107.07280}

%\bibitem[DS]{DS}
%Drinfeld V. G., Sokolov V. V., Lie algebras and equations of Korteweg-de Vries type. Journal of Soviet Mathematics, 1985, 30:2, 1975--2036

%\bibitem[F]{F}
%Freeman M.D., On the mass spectrum of affine Toda field theory. Phys. Lett. B261 (1991), 57--61

%\bibitem[FKM]{FKM}
% Freund P. G. O.,  Klassen T. R., Melzer E., $S$-matrices for perturbations of certain conformal field theories. Phys. Lett. B 229 (1989), no. 3, 243--247

%\bibitem[Ga]{Ga}
%Gantmacher F. R.,  \textit{The Theory of Matrices} Trans. from the Russian by K. A. Hirsch, vols. I and II. New York, Chelsea, 1959. 10+374 pp.; 9+276 pp.

\bibitem[GLO]{GLO}
Gerasimov A.A.,  Lebedev D.R., Oblezin S.V.,
On quantum $\fosp(1|2\ell)$-Toda chain; \url{https://arxiv.org/pdf/2012.08902}

\bibitem[GL]{GL}
Grozman P., Leites D., Defining relations associated with the principal
$\mathfrak{sl}(2)$-subalgebras. In: Dobrushin R.,
Minlos R., Shubin M. and Vershik A. (eds.) \textit{Contemporary
Mathematical Physics} (F.~A.~Berezin memorial volume), Amer. Math.
Soc. Transl. Ser. 2, vol. 175, Amer. Math. Soc., Providence, RI
(1996), 57--67; \url{https://arxiv.org/pdf/math-ph/0510013}.

\bibitem[GL]{GL}
Grozman P., Leites D., From supergravity to ballbearings. In:
J.~Wess, E.~Ivanov (eds.), \textit{Supersymmetries and quantum
symmetries}, (SQS'97, 22--26 July, 1997), Lecture Notes in Phys.,
524, (1999)  58--67

\bibitem[GL1]{GL1}
Grozman P., Leites D., Defining relations for classical Lie superalgebras with Cartan
matrix. Czech. J.~Phys., Vol. 51, 2001, no.~1,
1--22; \url{https://arxiv.org/pdf/hep-th/9702073}.


\bibitem[IvKr]{IvKr}
Ivanov E. A., Krivonos S. O., $U(1)$-supersymmetric extension of the Liouville equation. Lett. Math. Phys. 7 (1983), no. 6, 523--531.
%Ivanov E. A., Krivonos S. O., 
Errata: ``$U(1)$-supersymmetric extension of the Liouville equation''. Lett. Math. Phys. 8 (1984), no. 4, 345.

\bibitem[Ki]{Ki}
Kiselev A. V., Methods of geometry of differential equations in analysis of integrable models of field theory.
Journal of Mathematical Sciences (New York), (2006), 136:6, 4295--4377


%\bibitem[KV1]{KV1}
%Govind S. Krishnaswami, T. R. Vishnu, The idea of a~Lax pair --- Part I: Conserved quantities for a~dynamical system, Resonance, Vol. 25, No. 12  (2020) 1705--1720,

%\bibitem[KV2]{KV2}
%Govind S. Krishnaswami, T. R. Vishnu, The idea of a~Lax pair ---  Part II: Continuum Wave Equations. Resonance, Vol.  26, (2021) 257--274 (An introduction to Lax pairs and the zero curvature representation;  \url{https://arxiv.org/pdf/2004.05791})

\bibitem[Lg]{Lg}
Lang S.,  \textit{Differential and Riemannian manifolds}. Springer-Verlag. (1995).

\bibitem[Lsos]{Lsos}
Leites D.,  \textit{Supermanifold theory}. Karelia Branch of the USSR Acad. Sci., Petrozavodsk, 1983, 200 pp (in Russian), see [B1] in \url{https://staff.math.su.se/mleites/books.html}

\bibitem[L]{L}
Leites D.,  How to superize the Liouville equation. J. Nonlinear Math. Phys. 7 (2000), no. 3, 263--267;  \url{https://arxiv.org/pdf/hep-th/0007256}

\bibitem[L1]{L1}
Leites D.,  The Riemann tensor for nonholonomic manifolds. Homology,
Homotopy and Applications, vol 4 (2), (2002)  397--407;
\url{https://arxiv.org/pdf/math.RT/0202213}

\bibitem[L2]{L2}
Leites D.,  Two problems in the theory of differential
equations. Theor. and Mathem. Physics, Vol.  192(2) (2019) 271--283; \url{https://arxiv.org/pdf/1905.11754}

\bibitem[Lo]{Lo}
Leites D.,  On odd parameters in geometry. J. Lie Theory. Volume 33, no. ? (2023) 40 pp. (accepted); \url{http://arxiv.org/pdf/2210.17096}

%\bibitem[LL]{LL}
%Leites D., Lozhechnyk O., Inverses of Cartan matrices of Lie algebras and  Lie superalgebras. Linear Algebra  Appl.
%V. 583 (2019), 195--256; \url{https://arxiv.org/pdf/1905.12408}

%\bibitem[LL1]{LL1}
%Leites D., Lozhechnyk O., On a~hidden supersymmetry of cosmological billiards. \url{https://arxiv.org/pdf/2102.04192}

\bibitem[LSS]{LSS}
Leites D., Saveliev M., Serganova V., Embeddings of $\mathfrak{osp} (N|2)$ and completely integrable systems. In: M.~Markov,
V.~Man'ko (eds.) \textit{Proc. International Conf. Group-theoretical
Methods in Physics}, %Yurmala, May, 1985. Nauka, Moscow, 1986,
%377--394 MR 89h:17042 (English translation:
VNU Sci Press, (1987), 255--297.

%\bibitem[LeSe]{LeSe}
%Leites D., Serganova V., Solutions of the classical Yang--Baxter equation
%for simple Lie superalgebras.
%Theor. and Mathem. Phys. \textbf{58} (1984), no.~1, 16--24.

\bibitem[LS]{LS}
Leznov A. N., Saveliev M. V., Representation Theory and Integration
of Nonlinear Spherically Symmetric Equations to Gauge Theories. Commun. Math. Phys. 74,(1980) 111--118

\bibitem[LSb]{LSb}
Leznov A. N., Saveliev M. V., \textit{Group-theoretical methods for integration of nonlinear dynamical systems}. Translated from the Russian  and revised  by D. A. Leĭtes. Progress in Physics, 15. Birkh\"auser Verlag, Basel, (1992) xviii+290 pp.



\bibitem[LSL]{LSL}
Leznov A. N., Saveliev M. V., Leites D. A., Superalgebra $b(0,1)$ and explicit integration of the supersymmetric Liouville equation. Phys. Lett. B 96 (1980), no. 1--2, 97--99



\bibitem[Li]{Li}
\url{https://en.wikipedia.org/wiki/Liouville%27s_equation}

\bibitem[MaG]{MaG}
Manin Yu., \textit{Gauge field theory and complex geometry}. Second
edition. Springer-Verlag, Berlin, (1997) xii+346 pp.



%\bibitem[MOP]{MOP}
%Mikhailov A.V., Olshanetsky M.A., Perelomov A.M., Two-dimensional generalized Toda lattices. Commun. Math. Phys. 79 (1981), 473--488

\bibitem[O]{O}
Olshanetsky M. A., Supersymmetric two-dimensional Toda lattice. Comm. Math. Phys. 88 (1983), no.~ 1, 63--76

%\bibitem[OV]{OV}
%Onishchik A., Vinberg E., \textit{Lie groups and algebraic groups.}
%Springer-Verlag, Berlin, (1990) xx+328 pp.

%\bibitem[P]{P}
%Perelomov A. M., Remarks on the mass spectrum of two-dimensional Toda lattice of~ $E_8$ type. J. Nonlinear Math. Phys. 27 (2020), no. 1, 12--16.

\bibitem[Po]{Po}
Polyakov A.M., %Quantum Geometry of Bosonic Strings, Phys. Lett. B, (1981) V.103, 207--210; 
Quantum Geometry of Fermionic Strings, Phys. Lett. B, (1981) V.103, 211--213.

%\bibitem[Pr]{Pr}
%Prasolov V., \textit{Problems and Theorems in Linear Algebra}. Translations of Mathematical Monographs, Vol. 134, American Mathematical Society (1994),
%Chapter IV, 114--137 pp.

\bibitem[SaVe]{SaVe}
Saveliev, M. and Vershik, A.: Continuum analogues of contragradient Lie algebras, Comm. Math. Phys. 126 (1989), 367--378

%\bibitem[Sch]{Sch}
%Schmidtt D. M., Supersymmetry of affine Toda models as fermionic symmetry flows of the extended mKdV hierarchy. SIGMA (Symmetry Integrability Geom. Methods Appl.) 6 (2010), Paper 043, 25 pp.

%\bibitem[St]{St}
%Stekolshchik, R. \textit{Notes on Coxeter Transformations and the McKay Correspondence}, Springer Monographs in Mathematics, (2008), xx+239 pp.; \url{https://arxiv.org/pdf/math/0510216}

\bibitem[Top]{Top}
Toppan F., On anomalies in classical dynamical systems. J. of Nonlinear Mathem. Physics V. 8, (2001) no. 4, 518--533.

\bibitem[Va]{Va}
Vasiliev M. Higher Spin Gauge Theories: Star-Product and AdS Space. In: Shifman M. (ed.) \textit{The Many Faces of the Superworld}. World Scientific. (2000) 533--610; \url{https://arxiv.org/pdf/hep-th/9910096}

\bibitem[V]{V}
Vershik A. M. Mathematics of nonholonomicity. Appendix 3 in: Sergeev V., \textit{The Thermodynamic Approach to Markets}, (ca. 160 pp)
MPIMiS \url{http://www.mis.mpg.de}  preprint
81/2006; \url{https://arxiv.org/pdf/0803.3432}


%\bibitem[W]{W}
%Witten E., Notes on supermanifolds and integration. Pure and Applied Mathematics Quarterly (Special Issue on Super Riemann Surfaces and String Theory)
%V. 15 (2019) No. 1,  3 --56

%\bibitem[Z]{Z}
%Zamolodchikov A.B., Integrals of motion and $S$-matrix of the (scaled) $\protect{T=T_c}$ Ising model with magnetic field. Int. J. Mod. Phys., 4 (16) (1989), 4235--4248.

\bibitem[ZP]{ZP}
Zupnik B. M., Pak D. G., Superfield formulation of the simplest three-dimensional gauge theories and conformal supergravities, Theoret. and Math. Phys., 77:1 (1988), 1070--1076

\end{thebibliography}

\documentclass[12pt,a4paper]{amsart}

\usepackage{hyperref}

\usepackage{DLdef1}

\let\ssec\subsection

\renewcommand {\ssbegin}[2][*]
 {\refstepcounter{subsection}%
\if#1*
\addcontentsline{toc}{subsection}{\thesubsection.\hskip 1pc #2}%
\else
\addcontentsline{toc}{subsection}{\thesubsection.\hskip 1pc #2. #1}%
\fi
 \def \secno {\gdef \secno {}{\ssecfont
\thesubsection.\hskip 2ex}%
 }%
 \begin{#2}}

\renewcommand {\sssbegin}[2][*]
  {\refstepcounter{subsubsection}
\if#1*
\addcontentsline{toc}{subsubsection}{\thesubsubsection\hskip 1pc #2}%
\else
\addcontentsline{toc}{subsubsection}{\thesubsubsection\hskip 1pc #2. #1}
\fi
  \def \secno {\gdef \secno {}{\ssecfont \thesubsubsection\hskip 2ex}%
  }%
   \begin{#2}}

\renewcommand {\parbegin}[2][*]
  {\refstepcounter{paragraph}
\if#1*
\addcontentsline{toc}{paragraph}{\theparagraph.\hskip 1pc #2}%
\else
\addcontentsline{toc}{paragraph}{\theparagraph.\hskip 1pc #2. #1}
\fi
  \def \secno {\gdef \secno {}{\ssecfont \theparagraph.\hskip 2ex}%
  }%
   \begin{#2}}

\makeatletter
\newcommand\Subsubsection[2]{%
\refstepcounter{subsubsection}%
\addcontentsline{toc}{subsubsection}{\thesubsubsection.\hskip 1pc #1. #2}%
\par\vskip1ex \@plus .4ex \@minus .2ex\noindent\thesubsubsection. \normalfont{\bfseries #1} (#2).\hskip.7em plus .3em}
\makeatother

\begin{document}

\title{Superized Leznov-Saveliev equations as the zero-curvature condition on a reduced connection}

\author{Dimitry Leites}
\address{Department of Mathematics, Stockholm University, 
Stockholm, Sweden\\ 
dimleites@gmail.com}

\makeatletter
\@namedef{subjclassname@2020}{\textup{2020} Mathematics Subject Classification}
\makeatother

\keywords {Lie superalgebra, Toda lattice, reduced connection, Leznov-Saveliev equation}
 \subjclass[2020]{Primary  70S15, 81T13; Secondary 35Q70}

\begin{abstract} The equations of open 2-dimen\-sional Toda lattice  (TL) correspond to Leznov-Saveliev equations (LSE) interpreted as zero-curvature Yang-Mills equations  on the variety of $O(3)$-orbits on the Minkowski space when the gauge algebra is the image of $\fsl(2)$ under a principal embedding into  a simple finite-dimensi\-o\-nal Lie algebra $\mathfrak{g}(A)$ with Cartan matrix $A$. 

The known integrable super versions of TL equations correspond to matrices $A$ of two different types. I interpret the super LSE of one type 1 as zero-curvature equations for the \textit{reduced} connection on the non-integrable distribution on the supervariety of $OSp(1|2)$-orbits on the $N=1$-extended Minkow\-ski superspace; the Leznov-Saveliev method of solution is applicable only to  $\mathfrak{g}(A)$ finite-dimensional and admitting a superprincipal embedding $\mathfrak{osp}(1|2)\tto\mathfrak{g}(A)$. The simplest LSE1 is the super Liouville equation; it can be also interpreted in terms of the superstring action.

Olshanetsky introduced LSE2 --- another type of equations of super TL. 
Olshanetsky's equations, as well as LSE1 with infinite-dimensional $\mathfrak{g}(A)$, can be solved by the Inverse Scattering Method. To interpret these equations remains an open problem, except for the super Liouville equation --- the only case where these two types of LSE coincide.

I also review related less known and less popular mathematical constructions involved.

\end{abstract}


\maketitle

\markboth{\itshape 
Dimitry Leites} 
{{\itshape An interpretation of certain superized Leznov-Saveliev equations}}

\thispagestyle{empty}

\section{Introduction}

The zero curvature gauge field equations were superized in two ways yielding solvable equations. Their meaning was not, however, explained except for the simplest case of super Liouville equation whose interpretation is not related with gauge fields, but with a~superstring action, see \cite{Po}.

Here, I recall two approaches to solutions of zero-curvature gauge field equations. I review the subtleties of their superizations needed to understand what is going on, e.g., appearance of integrable forms, see their definition \eqref{intForm}. Vital in the \textbf{super} version of gauge field (a.k.a. Yang-Mills) equations is a~ \textbf{non-integrable distribution} on the supermanifold considered, such as Minkowski supermanifold; this was never explained before. The methods of solution in these approaches are different and the ranges of their applicability intersect in one case only.  


\ssec{From the Yang--Mills equations to the variety of $\text{O}(3)$-orbits in the Minkowski space and further to zero-curvature equations with the gauge algebra $\fsl(2)\subset \fg(A)$. Two methods to solve these equations}  In physics, a~\textit{gauge theory} is a~type of field theory in which  the dynamics of the system is invariant  under local transformations performed by Lie groups, or --- infinitesimally --- Lie algebras. 

The equations of gauge theories --- called \textit{Yang--Mills} equations --- are too complicated to be explicitly solved; solutions are found only under some additional restrictions.

\textbf{1)  The Leznov--Saveliev method}. Under additional spatial $\text{O}(3)$-symmetry, the 4-dimensional Minkowski space reduces to a~2-dimensional cylinder with 2 parameters: the radius $r$ of the $\text{O}(3)$-orbit (the sphere in $\Ree^3$) and the time $t$. In the simplest case of zero-curvature conditions on the gauge field, these equations  --- called \textit{Leznov--Saveliev} equations (LSE) --- can be solved; the method of solution requires the complexified gauge Lie algebra $\fg$ to be of finite dimension and admitting a principle embedding of $\fsl(2)$. The Leznov--Saveliev equations do not, however, consider abelian theories, they are applicable only to simple finite-dimensional gauge Lie algebras $\fg$, hence, of the form $\fg=\fg(A)$ with Cartan matrix $A$; so the LSE are of interest as integrable systems, but are hardly applicable to certain current models of particle physics with abelian summand in the gauge algebra. 

Explicitly, the \textit{Leznov-Saveliev equations} constitute the system \eqref{LS} of equations for a~vector-valued function $F=(F_1, \dots, F_n)$, where $n=\rk\fg$ and $\fg=\fg(A)$ is a~ complex  Lie algebra with Cartan matrix $A:=(A_{ij})_{i,j=1}^n$, each $F_i$ depends on $x$ and $y$ which are complexified linear combination of $r$ and $t$:
\be\label{LS}
\nfrac{\partial^2 F_i}{\partial_x\partial_y}=\exp\left(\sum_j A_{ij}F_j\right),\text{~~where $i=1,\dots , n$}.
\ee
Sometimes,  the term \textit{Leznov-Saveliev equations} is applied to the following system of equations for a~vector-valued function $F=(F_1, \dots, F_n)$, where each $F_i$ depends on $x$ and $y$:
\be\label{LSbis}
\nfrac{\partial^2 G_i}{\partial_x\partial_y}=\sum_j A_{ij}\exp(G_j),\text{~~where $i=1,\dots , n$}.
\ee

If the matrix $A$ is invertible (in the simplest case, $\fg(A)$ is finite-dimensional and simple),  then the system \eqref{LS}  is equivalent to the system~ \eqref{LSbis}: set $G=AF$, and hence $F=A^{-1}G$.

If $A$ is  the $n\times n$ Cartan matrix of a~simple finite-dimensional Lie algebra (over $\Cee$), then the general solution of  eq.~\eqref{LS} is a~vector depending on (the logarithm in) $2n$ polynomials in 2 arbitrary analytic functions each (one function depending on $x$, the other one on $y$) and their derivatives, see \cite{LS, LSb}.

\textbf{2)  The Inverse Scattering Method}. This powerful method was  discovered earlier than the LS-method; it is applicable not only to finite-dimensional gauge algebras $\fg(A)$ but also to algebras of infinite-dimension  (namely, affine Kac-Moody algebra a.k.a. \textit{double extension} of the loop algebra, see the last two Sections in \cite{BLS}), see, e.g., \cite{AC, AN}. This method allows one to solve versions of Leznov--Saveliev equations (sine-Gordon and sinh-Gordon equations in the simplest cases) to which the Leznov--Saveliev method is inapplicable: there are no principal embeddings of $\fsl(2)$ into the infinite-dimensional Lie algebras.

If $A$ is  the Cartan matrix of an infinite-dimensional Kac-Moody Lie algebra, the solution is obtained via ISM in terms of theta-functions, see \cite{AN}.

\ssec{Superization of Leznov-Saveliev equations via the  LS-approach} It seems natural to consider equations \eqref{LS} and \eqref{LSbis} where $A$ is any Cartan matrix of any finite-dimensional or Kac-Moody Lie superalgebra. However, to this day only two types of Cartan matrices were considered in order to get a solution of these equations, and only one of these two types of equations has an interpretation: it is given in this paper.

\sssec{Non-integrable distributions}   
Recall that a \textit{distribution} $\cD$ on the manifold or supermanifold $M$ is a subbundle of the tangent bundle $TM$; it is given by selecting a subspace $\cD_m$ of the tangent space $T_mM$ at every point $m\in M$; we assume that all spaces $\cD_m$ are of the same dimension $k$ and smoothly depend on $m$. The distribution $\cD$ is \textit{integrable} at a point $m\in M$ if there is a $k$-dimensional submanifold $N\subset M$ containing $m$ such that $T_mN=\cD_m$.
For integrability of the distribution, there is the following criterion true on supermanifolds as well.

\paragraph{The Frobenius Criterion for integrability of the distribution (\cite[Ch. VI]{Lg})} \textit{The distribution $\cD$ is integrable if and only if its sections form a Lie subalgebra of $\fvect(M)$ closed under the Lie bracket of vector fields}.

\sssec{$N$-extended Minkowski supermanifolds and non-integrable distributions} The advent of supersymmetry which undoubtedly is the language and tool for implementing  the Einstein's dream --- the future Grand Unified Theories of all fundamental forces, called SUSY GUTS --- brought new direction of research: solution of the  Yang--Mills  equations on supermanifolds. Observe that the Frobenius criterion for integrability of distributions is true on supermanifolds; for a~ proof, see \cite{Lsos}.

Observe that the $N$-extended Minkowski supermanifold $\cM_N$, where\footnote{This restriction is imposed ``for physical reasons", in order to eliminate particles of spin $>2$. Since 1990, M.~Vasiliev with co-authors and followers develops a theory which lacks this restriction, see one of the latest papers \cite{Va} and references therein.} $N=1, \dots, 8$, is not just a supermanifold: it is endowed with a non-integrable distribution. A precise shape of $\cM_N$ is not known for $N>1$; for a discussion, see \cite{MaG, GL}. However, it is clear from the ``twistor'' (see \cite{MaG, GL}) interpretation of the Minkowski supermanifold $\cM_N$ as the quotient space of a real form of the complex Lie supergroup $\SL(4|N)$, the Lie superalgebra of this real form $\cG$ consisting of the following supermatrices in the \lq\lq non-standard" format $2|N|2$ (here $\overline X$ denotes the matrix obtained from $X$ by element-wise complex conjugation and ${}^t$ is the sign of transposition):
\begin{equation}\label{M3}
\begin{pmatrix}
A&\overline R^t&U\\
Q&B&R\\
T&-\overline Q^t&-\overline A^t
\end{pmatrix},\;\;\begin{array}{l}\text{where  $Q\in\Mat_\Cee(2\times N)$,}\ \ \text{$T=\overline T^t$, $R\in\Mat_\Cee(2\times N)$,
$U=\overline U^t$,}\\
\text{$A\in\fgl(2;\Cee)$,
 $B\in\fu(N)$, $\tr B=\tr(A-\overline A^t)$}\end{array}
\end{equation}
so $\cM_N=\cG/\cP$, where the subsupergroup $\cP$ is generated by block non-strictly-upper-triangular supermatrices.

Therefore, the tangent space to any point of $\cM_N$ is isomorphic to the superspace spanned by block supermatrices $Q$, $\overline Q^t$, and $T$; the distribution of codimension $4|0$ spanned  by $Q$ and $\overline Q^t$ at the tangent space at every point is non-integrable since $[Q,\overline Q^t]=T$. Other models of $\cM_N$ for $N>1$ are described by smaller parabolic subgroup $\cP$, so a part of $B$ becomes a~ part of the tangent space at a~ point of $\cM_N$; the non-integrable distribution described above only becomes more complicated, see \cite{GL,BGLS}.

\sssec{LSE1 and reduced connections} The supervariety of $\text{OSp}(1|2)$-orbits in the $N=1$-Minkowski superspace is a~$2|2$ dimensional supermanifold, that inherits a~non-integrable distribution of codimension $2|0$, see \cite{LSL, LSS}.
The described in \cite{LSS} method  is applied to the zero curvature equations for a~\textbf{reduced connection}, as explained further on; at the time \cite{LSS} was written the notion of reduced connection was unknown to the authors who performed superization of LSE guided  only by their gut feeling. The zero curvature equations for a~reduced connection on the supervariety of  these orbits --- \textit{superized  Leznov--Saveliev equations} or LSE1 --- can be solved if there is  a~ \textit{superprincipal} embedding $i:\fosp(1|2)\tto\fg$ into the gauge Lie superalgebra~ $\fg$. (Recall that the embedding $i$ is called \textit{superprincipal} if the adjoint representation of $i(\fosp(1|2))$ in $\fg$ splits into the smallest possible number  (equal to the rank of $\fg$) of irreducible modules.) The simple Lie superalgebras admitting a superprincipal embedding into them must be  
\begin{equation}\label{lsA}
\begin{minipage}[c]{14cm} \textbf{finite-dimensional, have Cartan matrix with\\ only 0s and 1s on its main diagonal}.\end{minipage}
\end{equation}

Observe that some authors (e..g., \cite{A1, A2}) consider Cartan matrices $A$ that do not allow superprincipal embeddings to $\fg(A)$, although   $A$ satisfies condition~\eqref{lsA}; it is unclear if the corresponding super LSE1 are integrable  with an explicit answer, as in \cite{AN}, not as ``some formal series".

Observe that the simplest case of these superized LSE1 is the super Liouville equation which appeared in Polyakov's paper \cite{Po} from a totally different direction, not related with any gauge equation, but as a~ result of variation of the action in the description of the simplest fermionic string. 

\ssec{Olshanetsky's approach: superization LSE2} M.~Olshanetsky suggested a~ completely different approach to superizing TL equations, see \cite{O}.  Olshanetsky solved these LSE2 by means of the Inverse Scattering Method. Interestingly, the ISM is applicable to  infinite-dimensional gauge Lie superalgebra, as in periodic TL (provided it is ``not too big": namely be  not of exponential growth, as hyperbolic algebras or almost affine superalgebras as described in \cite{CCLL}), it should be
\begin{equation}\label{OlsA}
\begin{minipage}[c]{12cm} \textbf{not necessarily finite-dimensional, can be $\Zee$-graded of polynomial growth, and its Cartan matrix should have only 2s and 1s on its main diagonal}.\end{minipage}
\end{equation}
Observe that such Lie superalgebras have various properties almost identical to those of finite-dimensional simple Lie algebras --- unlike properties of Lie superalgebras with Cartan matrices of different shape or without Cartan matrix at all.


Observe that the equations suggested by Olshanetsky are more complicated than eq.~\eqref{SLS}; they appear from variation of a (seemingly \textit{ad hoc}) action functional. To  interpret these LSE2 remains an open problem unless $A=(1)$ in which case LSE2=LSE1, and there are two interpretations: in \cite{Po} and in this paper.

\section{The gauge fields a.k.a. connections} Following  \cite{MaG} and \cite{Lsos, L2} recall that there are two equivalent definitions of connections, which we consider directly in the ``super'' setting. Let $M$ be a~(super)manifold, $\cF$ the algebra of functions, $\fvect:=\fder(\cF)$ the Lie algebra of vector fields on $M$, and  $\Omega^1:=\Pi(\Hom_\cF(\fvect, \cF))$, where $\Pi$ is the reversal of parity functor, the space of differential 1-forms.
Let $\fg$ be a~Lie (super)algebra, let $\rho$ be a~representation of $\fg$ in the (super)space~ $V$. Let $T(V):=\cF\otimes V$ be the $\cF$-module of tensor fields with fiber $V$, so $\fg$ can be considered as a~Lie sub(super)algebra of $\End_\cF(T(V))\simeq\End(V)$.

1) Classically, a~\textit{connection} (or a~\textit{gauge field}) on $M$ in the $(\cF,\fvect)$-bimodule $T(V)$ is an \textbf{even} map
\be\label{conn}
\nabla:\fvect\times T(V)\tto T(V)
\ee
which is $\cF$-linear in the first argument, additive in the second argument and satisfies the following version of the Leibniz rule, where $\nabla_D(t):=\nabla(D, t)$
\[
\nabla_D(ft)=D(f)v+(-1)^{p(f)p(D)}f\nabla_D(t)\text{~~for any $f\in\cF$, $t\in T(V)$, $D\in\fvect$}.
\]
The operator $\nabla_D:=D+\rho(X)\in \End_\cF(T(V))\simeq\End(V)$, where the  summand $D$ is the standard shorthand for the diagonal  operator $D\otimes \id_V$ and $X\in \fg$, is called the \textit{covariant derivative} along~ $D$.

The \textit{curvature} of the connection $\nabla$ is defined as the tensor
\[
R(D_1, D_2):=\nabla_{D_1}\nabla_{D_2}-(-1)^{p(D_1)p(D_2)}\nabla_{D_2}\nabla_{D_1} -\nabla_{[D_1, D_2]}.
\]

2) Dualizing the above definition, and changing parity we get the following, more modern, definition allowing us to explore higher differential forms and --- on supermanifolds --- integrable forms, i.e., the form one can integrate, see \cite{BL}. 
Unlike \cite{ZP},  we will not use this definition; we give it here for completeness of the review part of this paper.

The \textbf{odd} operator $\nabla=d+\alpha$ is called a~\textit{connection} (or a~\textit{gauge field}) on $M$; here $d$ is a~shorthand for the diagonal  operator $d\otimes \id_V$, and
\[
\alpha\in\Omega^1\otimes \fg\subset \Omega^1\otimes_{\cF} \End_{\cF}(T(V))=
\Omega^1\otimes\End(V),
\]
denotes both the~$\fg$-valued differential 1-form  on  $M$ and the operator of left multiplication by it. Observe that on manifolds, any 1-form $\alpha$ is odd, so the sum $d+\alpha$ is automatically homogeneous, whereas on supermanifolds we have to require $\alpha\in(\Omega^1)_\od$. In the cases where a~naturally selected $\alpha$ is even, as on the supermanifolds with a~pericontact form $d\tau+\sum \pi_iq_i$ with odd ``Time'' $\tau$ and ghosts $\pi_i$, we have to consider connections with odd parameters, see \cite{L2,  Lo}.

The form $F_\nabla:=\frac12[\nabla, \nabla]$, where the commutator is taken in $\fg\otimes \Omega$ for $\Omega:=\oplus \Omega^i$, is called the \textit{curvature} of $\nabla$.

Let $(V_i, \nabla_i)$ for $i=1,2$ be a shorthand for modules with connections $(T(V_i),\nabla_i)$.
Define their tensor product and the module of homomorphisms  by setting for  ${(V_1\otimes_{\cF} V_2, \nabla_1\otimes \nabla_2)}$ and $(\Hom_{\cF}(V_1, V_2), \ \nabla^{\Hom(V_1, V_2)})$ 
\[
\begin{array}{l}
(\nabla_1\otimes \nabla_2)(v_1\otimes v_2):=\nabla_1(v_1)\otimes v_2+(-1)^{p(v_1)}(T\otimes \id_{V_2})(v_1\otimes \nabla_2(v_2)),\\
\nabla^{\Hom(V_1, V_2)}(\varphi)(v_1):=\nabla_2(\varphi(v_1))-(-1)^{p(\varphi)}(\id_{V_1} \otimes \varphi)(\nabla_1(v_1))
\end{array}
\]
for any $v_i\in V_i$ and $\varphi\in \Hom_{\cF}(T(V_1), T(V_2)))$, where
\[
T: \Omega^1\otimes V\simeq  V\otimes\Omega^1
\]
is the twisting isomorphism. In particular, set $\nabla^{\End(V)}:=\nabla^{\Hom(V, V)}$.

\ssec{The Yang--Mills equation} The \textit{Yang-Mills equation} is
\be\label{YM}
\nabla^{\End(V)}((F_\nabla)^*)=0.
\ee

Observe that the ``zero curvature condition" $F_\nabla=0$ is invariant under any changes of coordinates
and does not depend on the choice of one of the two definitions of the connection. The ``zero curvature condition" is the simplest version of YME.

On the $n$-dimensional manifold $M$, we have
\be\label{dual}
(\Omega^i)^*\simeq\Omega^{n-i},
\ee
where the dualization is considered over the algebra $\cF=\Omega^0$ of functions, not over the ground field. Although the isomorphism \eqref{dual} is correctly defined only for forms with compact support or whatever other type of forms for which the integral is well defined, we will use it ``formally'', recklessly ignoring the questions of convergence. 

Similarly, on the supermanifold $\cM$, the dual to the space of \textit{differential $i$-forms} defined to be elements of the space $\Omega^i:=T(E^i(V))$, where $E^i$ is the operator of raising to the $i$th exterior power and the cotangent space $V$ at the point of $\cM$ is the tautological $\fgl(V)=\fgl(\dim(\cM))$-module, is the space of \textit{integrable $(-i)$-forms}
\begin{equation}\label{intForm}
\Sigma_{-i}(\cM):=(\Omega^i(\cM))^*\otimes_\cF \Vol(\cM)\simeq T(E^i(V^*)\otimes \tr),
\end{equation}
where $\Vol(\cM)$ is the $\cF$-module of volume forms (a.k.a.  Berezinian), and $\tr$ denotes both the (super)trace on $\fgl(\dim\cM)$ and the 1-dimensional (even) module it defines. There are two complexes (series of mappings such that the product of two consecutive mappings is equal to 0)
\be\label{d}
d\colon \Omega^i \tto\Omega^{i+1} ; \ \
d^*\colon \Sigma_{-i-1} \tto \Sigma_{-i}.
\ee
On supermanifolds, $(F_\nabla)^*$ in the Yang-Mills equation \eqref{YM} is an integrable form,  see \cite{ZP}.

\ssec{The Leznov-Saveliev equations} Here is how Yang-Mills equations turn into LSE.

Recall that the Minkowski space $M$ is, locally, $\Ree^{3,1}$ on which the orthogonal groups $\text{O}(3)$ naturally acts rotating its $\Ree^3$ part. The variety of orbits are 2-dimensional ``cylinders'', each being ``$\text{sphere $S^2$}\times \text{Time}$", with coordinates being the radius of the sphere $r$ and the Time parameter~ $t$. Clearly, the Poincar\'e group preserving the Lorentz metric contains $\text{O}(3)$. Now, let us complexify everything and consider the linear approximation, i.e., pass from Lie groups to their Lie algebras.
Consider a connection whose gauge Lie algebra is the image of the principal embedding $\fsl(2)\tto \fg$  into the Lie algebra of the gauge group $G$. The LSE are the  zero curvature conditions on the variety of $\text{O}(3)$-orbits.

Recall that an embedding $i:\fsl(2)\tto\fg$, where $\fg$ is a~simple finite-dimensional Lie algebra,  is called \textit{principal}, if $\fg$ splits into $\rk \fg$ irreducible $i(\fsl(2))$-submodules. Every simple finite-dimensional Lie algebra $\fg$ admits a~principal embedding, all these embeddings are isomorphic under the Lie group of automorphisms  of $\fg$.
After complexification and passage to the ``light cone coordinates'', i.e., instead of d'Alembertian in $r$ and $t$ we consider the Laplacian in $x$ and $y$, we arrive at the LSE \eqref{LS}.

The 1-dimensional Toda lattice (TL) is related to its version described by the  Leznov-Saveliev equations as explained on the first page of \cite{Ki}; both TL and LSE depend on 2 variables, but --- for reasons unknown --- only LSE is called \lq\lq 2-dimensional".

\sssec{The simplest example}\label{ssSimEx} Consider the following basis of $\fg=\fsl(2)$:
\[
X^-:=\begin{pmatrix} 0&0\\
1&0\end{pmatrix},\ \ H:=\begin{pmatrix} 1&0\\
0&-1\end{pmatrix},\ \ X^+:=\begin{pmatrix} 0&1\\
0&0\end{pmatrix}.
\]

In terms of the first definition of connection, Leznov and Saveliev (see \cite{LS, LSb}) considered the following $\fsl(2)$-valued connection
\be\label{Zcurv}
\begin{array}{l}
\nabla_{\partial_x}:=\partial_x+aH+bX^+,\\
\nabla_{\partial_y}:=\partial_y+AH+BX^-.
\end{array}
\ee

The zero-curvature condition takes the form
\[
[\partial_x+aH+bX^+, \ \
\partial_y+AH+BX^-]=0,
\]
which is the same as (here $f_x:=\frac{\partial f}{\partial x}$ and $f_y:=\frac{\partial f}{\partial y}$)
\[
A_xH+B_xX^--a_yH-b_yX^+-2aBX^--2bAX^++bBH=0
\]
equivalent to the system of equations
\[
\begin{array}{l}
A_x-a_y=-bB,\\
B_x=2aB,\\
b_y=-2bA,\\
\end{array}
\Longleftrightarrow
\begin{array}{l}
A_x-a_y=-bB,\\
(\ln B)_x=2a,\\
(\ln b)_y=-2A,\\
\end{array}
\]
wherefrom we derive the Liouville equation \eqref{LE}
\[
(\ln b)_{xy}+(\ln B)_{xy}=2bB \Longleftrightarrow F_{xy} =2\exp(F), \text{~~where $F=bB$}.
\]

\sssec{General LSE and zero-curvature equations} Let $\fg=\oplus \fg_i$ be any $\Zee$-graded Lie (super)algebra, $\fg_0=\Span(H_i)_{i\in I}$ and
$\fg_{\pm 1}=\Span(X_j^{\pm})_{j\in J}$. Let $I=J$ which is true in many cases (e.g., if $\fg$ has a Cartan matrix, and can be further generalized to Lie (super)algebras with ``continuum root system", see \cite{SaVe}). Leznov and Saveliev suggested to consider the following connections on the variety of $O(3)$-orbits in the Minkowski space:
\be\label{ZcurvGen}
\begin{array}{l}
\nabla_{\partial_x}:=\partial_x+\sum(a_iH_i+b_iX_i^+),\\
\nabla_{\partial_y}:=\partial_y+\sum(A_iH_i+B_iX_i^-).
\end{array}
\ee

The zero-curvature equation in this case leads to LSE only if $\fg$ has a Cartan matrix; this LSE can be explicitly  (resp. implicitly) solved only if $\fg$ is finite-dimensional (resp. of polynomial growth in Chevalley generators), see~ \cite{LS, LSb}.

\sssec{The Liouville equation}
The \textit{Liouville equation} can be written in various forms  (see \cite{Li}), in particular, as the simplest of Leznov--Saveliev equations:
\be\label{LE}
\nfrac{\partial^2 F}{\partial_x\partial_y}=\exp(2F).
\ee
Its general solution depends on two arbitrary analytic functions $f$ and $g$:
\be\label{2an}
F(x,y)=\frac12\ln\nfrac{f'(x)g'(y)}{(f(x)+g(y))^2}.
\ee
This fact is obvious if we somehow find one solution (for example, $f(x)=x$ and $g(y)=y$) and observe the conformal symmetry of the Liouville equation under the  action of the Lie algebra $\fwitt(x)\oplus\fwitt(y)$, where $\fwitt(x):=\fder\Cee[x^{-1}, x]$, called the \textit{Witt algebra}, is the Lie algebra of vector fields on the circle with coefficients expandable in Laurent polynomials, i.e., $x:=\exp(i\varphi)$, where $\varphi$ is the angle parameter on the circle, see \cite{L, IvKr}. (Observe that it is absolutely not obvious that the Lie algebra $\fvir(x)\oplus\fvir(y)$, where $\fvir(x)$ is the  \textit{Viraroso algebra}, the only non-trivial central extension of $\fwitt(x)$, also acts on the Liouville equation, but this is true, see \cite{Top}. Physicists call this phenomenon a~\textit{classical anomaly}.)

\ssec{The superized Leznov-Saveliev equations LSS1 (\cite{LSS})} A \textit{$2|2$-dimensio\-nal Toda lattice} is a~system of equations for the vector-function $F$ depending on two even variables $x, y$ and two odd (anti-commuting) indeterminates $\xi, \eta$ (they are not variables, since can only take value $0$)
\be\label{SLS}
D^+D^-F_i=\exp(\sum_j A_{ij}F_j) \text{~~for $D^+:=\partial_{\xi}+\xi\partial_{x}$ and $D^-:=\partial_{\eta}+\eta\partial_{y}$}
\ee
or its version (equivalent to eq.~\eqref{SLS} if $A$ is invertible: set $G=AF$, and hence $F=A^{-1}G$)
\be\label{SLSbis}
\frac{\partial^2 G_i}{\partial_x\partial_y}=\sum_j A_{ij}\exp(G_j),\text{~~where $i=1,\dots , n$}.
\ee


\sssec{The simplest example} Recall, see \cite{V}, that the \textit{reduced connection} on the (super)manifold with a non-integrable distribution is determined (\`a la the first definition of connection) on the fibers of the distribution, cf. with definition \eqref{conn}:
\be\label{redconn}
\nabla:\cD\times V\tto V
\ee
which is $\cF$-linear in the first argument, additive in the second argument and satisfies the following version of the Leibniz rule, where $\nabla_D(v):=\nabla(D, v)$
\[
\nabla_D(fv)=D(f)v+(-1)^{p(f)p(D)}f\nabla_D(v)\text{~~for any $f\in\cF$, $v\in V$, $D\in\fvect$}.
\]
Observe that the supervariety of orbits of the $\text{OSp}(1|2)$-action on Minkowski  supermanifold is of superdimension $2|2$, this supervariety inherits a non-integrable distribution induced by the distribution given on the Minkowski supermanifold, and the connection on the supervariety of orbits is determined by a reduced connection on the $(0|2)$-dimensional fibers of the distribution.

Consider the following basis of $\fosp(1|2)$:
\[
\begin{array}{l}
X^-:=\begin{pmatrix} 0&0&0\\
0&0&0\\
1&0&0\end{pmatrix},\ \ H:=\begin{pmatrix} 1&0&0\\
0&0&0\\
0&0&-1\end{pmatrix},\ \ X^+:=\begin{pmatrix} 0&0&1\\
0&0&0\\
0&0&0\end{pmatrix},\\
\hskip2cm \delta^-:=\begin{pmatrix} 0&0&0\\
1&0&0\\
0&1&0
\end{pmatrix},\ \ \delta^+:=\begin{pmatrix} 0&1&0\\
0&0&-1\\
0&0&0
\end{pmatrix}.
\end{array}
\]

Leznov,  Saveliev and Leites, see \cite{LSL} further developed in \cite{LSS}, considered the following $\fosp(1|2)$-valued \textit{reduced connection}  inherited from the $N=1$-Minkowski superspace:
\be\label{0curv}
\begin{array}{l}
\nabla_{+}:=D^++\alpha H+a \delta^+,\\
\nabla_{-}:=D^-+\beta H+ b\delta^-.
\end{array}
\ee
Let $\alpha, \beta\in\cC_\od$ and $a,b \in\cC_\ev$ for an auxiliary supercommutative superalgebra $\cC$ which will automatically disappear at the final stage.

After simplification similar to the one performed in Subsection \ref{ssSimEx}, the \textbf{reduced zero-curvature condition} $[\nabla_{+},
\nabla_{-}]=0$ for the \textit{reduced connection}  takes the form
equivalent to the system of equations
\[
\begin{array}{l}
D^+(\beta)+D^-(\alpha)=-ab,\\
D^+(b)=-\alpha b,\\
D^-(a)=a\beta,\\
\end{array}
\Longleftrightarrow 
\begin{array}{l}
D^+(\beta)+D^-(\alpha)=-ab,\\
D^+(\ln b)=-\alpha,\\
D^-(\ln a)=\beta,\\
\end{array}
\]
wherefrom we derive the \textit{super Liouville equation} \eqref{SLS} or \eqref{SLSbis} for just one function $F$ and $(A_{ij})_{i,j=1}^1=(1)$:
\be\label{SupLiou}
\begin{array}{l}
D^+D^-(\ln a)+D^-D^+(\ln b)=ab \Longleftrightarrow 
D^+D^-(F) =\exp(F), \text{~~where $F=ab$}.
\end{array}
\ee

Observe that the zero-curvature condition for the non-reduced connection \eqref{conn} would require two more conditions
\[
\begin{array}{l}
{}[\nabla_{+}, \nabla_{+}]=0,\\
{}[\nabla_{-}, \nabla_{-}]=0,\\
\end{array}
\]
which are never true since $(D^+)^2=\frac12[\nabla_{+}, \nabla_{+}]=\partial_x$ and $(D^-)^2=\frac12[\nabla_{-}, \nabla_{-}]=\partial_y$. For  analogs of the curvature tensors of non-integrable distributions on manifolds and supermanifolds, see \cite{L1, BGLS}.

The super Liouville equation was first solved in \cite{LSL} in components. For a~ conceptual solution in terms of superfields based on the supersymmetry relative $\fk^L(1|1)\oplus \fk^L(1|1)$, see \cite{IvKr}; to superize \cite{Top}, i.e., compute the value of the central charge with which the Neveu-Schwarz Lie superalgebra $\fns$ --- the non-trivial central extension of the Lie superalgebra $\fk^L(1|1)$ of contact vector fields on the $1|1$-dimensional superstring --- acts on the super LSE is an \textbf{open problem}, see~  \cite{L}.

The equation \eqref{SupLiou} is the zero-curvature condition for the reduced connection  on the supervariety of $\OSp(1|2)$-orbits acting on the simplest (complexified) $N=1$ Minkowski superspace  $\cM$, i.e.,  Yang--Mills equations invariant with respect to $\OSp(1|2)$ acting on (the space of functions, differential and integrable forms on) $\cM$, and on $\fg$ via a~superprincipal embedding $\fosp(1|2)\tto\fg$. 
Not every simple finite-dimensional Lie superalgebra $\fg$ admits superprincipal embeddings of $\fosp(1|2)$; but if such an embedding is possible, all superprincipal  embeddings are isomorphic under the group of automorphisms  of $\fg$. The simple finite-dimensional Lie superalgebras that do admit a~superprincipal embedding are $\fsl(n|n\pm1)$, and $\fosp(2n\pm1|2n)$, $\fosp(2n|2n)$, $\fosp(2n+2|2n)$, and  $\fosp_a(4|2)$, see \cite{LSS}.

\subsection*{Acknowledgements} I am thankful to 
A.~Lebedev for his clarifications and to the grant AD 065 NYUAD for financial support. 

\end{document}